It takes 18 employees to complete a job in 95 days. I want the job completed in 72 days, how many employees do I need?
Answer by ptaylor(2198)

Let x=number of employees it takes to complete the job in 72 days
It takes 18*95=1710 employee-days to do the job
In other words, 1 employee can complete the job in 1710/1=1710 days
2 employees in 1710/2
-
-
18 employees in 1710/18=95 days
x employees in 1710/x=72 days
x=23 3/4 employees. Since we can't have 3/4 of an employee, the answer has to be 24
CK
23.75*72=1710
Hope this helps---ptaylor